OpenStudy (yb1996):

What is the indefinite integral of x*sqrt((1-x)/(1+x)) ?

OpenStudy (freckles):

\[\int\limits x \sqrt{\frac{1-x}{1+x}} dx ?\]

OpenStudy (yb1996):

Yes

OpenStudy (freckles):

ok i think I might have an idea i haven't fully explored it

OpenStudy (anonymous):

put x=cos theta dx=-sin theta d theta

OpenStudy (freckles):

but try multiplying the inside by (1-x)/(1-x)

OpenStudy (freckles):

the inside of the square root

OpenStudy (freckles):

then we are gong to try a trig sub

OpenStudy (freckles):

\[\int\limits \frac{ x(1-x)}{\sqrt{1-x^2}} dx \\ \text{ use } x= \sin(\theta)\]

OpenStudy (yb1996):

Oh, that makes a lot of sense, I'll try that

OpenStudy (fibonaccichick666):

I like that freckles.

OpenStudy (freckles):

and thanks :)

OpenStudy (anonymous):

\[ \int\limits \cos \theta \sqrt{\frac{ 1-\cos \theta }{ 1+\cos \theta }}*-\sin \theta d \theta\] \[=\int\limits -\sin \theta \cos \theta \sqrt{\frac{ 2\sin ^2\frac{ \theta }{ 2 } }{ 2\cos ^2\frac{ \theta }{ 2 } }}d \theta \] \[=-\int\limits 2\sin \frac{ \theta }{ 2 }\cos \frac{ \theta }{ 2 }\frac{ \sin \frac{ \theta }{ 2 } }{ \cos \frac{ \theta }{ 2 } } \cos \theta d \theta \] \[=-\int\limits 2 \sin ^2\frac{ \theta }{ 2 }\cos \theta d \theta \] \[=-\int\limits \left( 1-\cos \theta \right)\cos \theta d \theta \] \[=-\int\limits \cos \theta d \theta+\frac{ 1 }{ 2 }\int\limits 2\cos ^2 \theta d \theta \] =?

OpenStudy (yb1996):

Completing what surjithayer did, I would get: \[\frac{1}{2}\left(\frac{1}{2}\left(2θ+\sin \left(2θ\right)\right)\right)-\sin \left(θ\right)+C\]

OpenStudy (yb1996):

This is what I got so far using freckles method, can someone check if I did it right?

OpenStudy (freckles):

your last line of your work in that pdf bothers me

OpenStudy (freckles):

everything else in that pdf is good til that last line

OpenStudy (freckles):

\[-\cos(\theta)-\frac{\theta}{2}+\frac{1}{4} \sin(2 \theta)+C \\ \text{ or } \\ -\cos(\theta)-\frac{\theta}{2}+\frac{1}{4} 2 \sin(\theta) \cos(\theta)+C \\ \text{ now recall } x=\sin(\theta) \\ \text{ so } \cos(\theta)=\sqrt{1-x^2} \\ \text{ inputtin these should give } \\ -\sqrt{1-x^2}-\frac{\arcsin(x)}{2}+\frac{1}{2}x \sqrt{1-x^2}+C\]

OpenStudy (yb1996):

Ahhh, ok, I was having some trouble putting x = sin(theta) back into the equation, but the way you did it makes sense

OpenStudy (freckles):

following surjithayer's work... \[- \int\limits \cos(\theta) d \theta+\frac{1}{2} \int\limits 2 \cos^2(\theta) d \theta \\ - \int\limits \cos(\theta) d \theta +\frac{1}{2} \int\limits (1+ \cos(2 \theta)) d \theta \\ - \sin(\theta)+\int\limits \frac{1}{2}d \theta+\frac{1}{2} \int\limits \cos(2 \theta) d \theta \\ -\sin(\theta)+\frac{1}{2} \theta+\frac{1}{4} \sin(2 \theta)+C \\ -\sin(\theta)+\frac{1}{2} \theta+\frac{1}{4} 2 \sin(\theta) \cos(\theta)+C \\ -\sin(\theta)+\frac{\theta}{2}+\frac{1}{2} \sin(\theta) \cos(\theta)+C \\ \text{ where he chose } x=\cos(\theta) \\ \text{ so you would have } \sqrt{1-x^2}=\sin(\theta) \text{ here } \\ -\sqrt{1-x^2}+\frac{\arccos(x)}{2}+\frac{1}{2} \sqrt{1-x^2} x +C\] you should see these as the answer since arcsin(x)=-pi/2-arccos(x)

OpenStudy (freckles):

same answer*

OpenStudy (freckles):

if you aren't sure how I'm getting the other trig function in terms of x given the substitution you can always draw a right triangle

OpenStudy (yb1996):

Awesome, thank you so much for your help!

OpenStudy (yb1996):

I understand your reasoning completely, it was just that I was blanking on the trig identity sin(2x) = 2sin(x)cos(x)

OpenStudy (freckles):

for example \[\text{ if } x=\sin(\theta) \\ \text{ or if } \sin(\theta)=\frac{x}{1}\] we have:  using the Pythagorean theorem we should see we have to solve the following equation: \[adj^2+x^2=1^2 \\ adj^2+x^2=1 \\ adj^2=1-x^2 \\ adj=\sqrt{1-x^2}\]